OpenStudy (anonymous):

Given tan theta=4/3 and sin theta=-4/5, find cos theta. A.1/3 B.-1/3 C.3/5 D.-3/5 E. -SQRT 3/5

OpenStudy (souvik):

\[\tan \theta =\frac{ \sin \theta }{ \cos \theta }\] so \[\cos \theta=\frac{ \sin \theta }{ \tan \theta}\]

OpenStudy (anonymous):

-4/5 ------ 4/3?

OpenStudy (souvik):

right...

OpenStudy (anonymous):

thank you!

OpenStudy (souvik):

you are welcome........

OpenStudy (anonymous):

= -3/5 :D

OpenStudy (souvik):

correct!

OpenStudy (anonymous):

can you help me with one more pls?

OpenStudy (anonymous):

Which expression is equivalent to tan(-x)cos2x? A. -sin xcos x B. sin xcos x C. tan x + sin2x D. tan x − sin2x E. -tan x − sin2x

OpenStudy (souvik):

tan(-x)=-tanx and cos2x=cos^2x-sin^2x

OpenStudy (anonymous):

i dont understand this one

OpenStudy (anonymous):

to my understanding then e is the only possible answer...

OpenStudy (souvik):

your question is tan(-x)*cos2x ! =-tanx[cos^2x-sin^2x]

OpenStudy (anonymous):

but thats not an option